OpenStudy (anonymous):

A greenhouse that specializes in growing geraniums is divided into sections. The number of geranium pots in each section depends on the number of sprinklers in that section. There are x(7x − 2) pots of red geraniums and (x + 3)2 pots of pink geraniums in a section with x sprinklers. Which simplified expression represents the total number of red and pink geranium pots in a section with x sprinklers? 8x2 + 4x + 9 8x2 + 6x + 9 8x2 + 6x + 7 8x4 + 4x2 + 9

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

anything

OpenStudy (welshfella):

first expand both functions x(7x - 2) = x*7x - 2*x = ? can you finish this off?

OpenStudy (anonymous):

8 x^2+4 x+9

OpenStudy (welshfella):

yep

OpenStudy (welshfella):

7x^2 - 2x + x^2 + 6x + 9 = 8x^2 + 4x + 9

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

The box plots below show the lifespan, in months, of laptop batteries manufactured by four companies: box plot labeled Company A shows minimum at 19, first quartile at 19.5, median at 23, third quartile at 23.5, and maximum at 24. Company B shows minimum at 21, first quartile at 22, median at 24, third quartile at 25, and maximum at 26. Company C shows minimum at 20, first quartile at 21, median at 22, third quartile at 24, and maximum at 28. Company D shows minimum at 19, first quartile at 20, median at 22.5, third quartile at 23.5, and maximum at 27. Based on the data, which company's batteries have the highest median lifespan? Company A Company B Company C Company D

OpenStudy (welshfella):

hmm. the median is the line inside the box right? I'm pretty rusty at this stuff

OpenStudy (welshfella):

If so its company B